Q1: A particle moves from x = 3.0 m to x = 9.0 m in 3.4 s, reaching a final speed of 3.00 m/s once it reaches x = 9.0 m. What is the (uniform) acceleration during this interval?
1
Expert's answer
2020-09-22T15:36:09-0400

The equation of the particle movement with the constant acceleration is:


"\Delta x = v_0t + \dfrac{at^2}{2}"

where "\Delta x = 9m-3m = 6m" is the trevelled distance, "v_0" is initial speed, "t = 3.4s" is the time and "a" is the acceleration.

The equation for velocity is:


"v = v_0 + at"

where "v = 3 m\/s" is the final velocity of the particle.

Solving this two equations simultaneously, obtain:


"v_0 = v - at\\
\Delta x = (v-at)t + \dfrac{at^2}{2} = vt - at^2 + \dfrac{at^2}{2} = vt-\dfrac{at^2}{2}"

Expressing "a", get:


"a = \dfrac{2v}{t} - \dfrac{2\Delta x}{t^2} = \dfrac{2\cdot 3}{3.4} - \dfrac{2\cdot 6}{3.4^2} \approx 0.73m/s^2"

Answer. 0.73 m/s^2.